OpenStudy (anonymous):

3 friends in a business agree to share the profits in a ratio 3:8:4 how much does each partner recieve from a total profit of 45 000

OpenStudy (anonymous):

sum of ratios = 3+8+4 = 15 so first one gets 3/15 * 45000 = 3 * 3000 = 9000 second one gets 8/15 * 45000 = 8 * 3000 = 24000 and third one gets 4/15 * 45000 = 4 * 3000 = 12000